I have a hypotnuse of 20 and I have an adjacent of 12.The opposite is unknown.I know I have to use cos, but I'm still unsure of how to solve for this.
Found 2 solutions by Fombitz, rfer:
Answer by Fombitz(32388)

What are you solving for?

Answer by rfer(16322)

c^2-a^2=b^2
400-144=256
sq rt 256=16